OpenStudy (anonymous):

what is the value of the y variable in the solution to the following system of equations? 3x-6y=12 and -2x+3y=6

OpenStudy (anonymous):

multiply the second equation with 2 then solve y will be canceled you get x then sub x in any equation you will get 'y'

OpenStudy (johnweldon1993):

Either that or multiply the first equation by 2 and the 2nd by 3....this will cancel the 'x' variable and let you solve for 'y' directly... 2(3x-6y=12) so this becomes 6x - 12y = 24 and 3(-2x+3y=6) so this becomes -6x + 9y = 18 So now we combine our 2 equations 6x - 12y = 24 +(-6x + 9y = 18) ----------------- 0 -3y = 42 so we have -3y = 42 Divide both sides by -3 to solve for 'y' -3y = 42 --------- -3 -3 y = 42/-3 y = ???

OpenStudy (anonymous):

so y is -14?

OpenStudy (johnweldon1993):

That would be correct! :)

OpenStudy (anonymous):

Thank you :)

OpenStudy (johnweldon1993):

Anytime :)